a)


Transistors of a certain type are specified to have β values in the range 50


to 150. Find the range of their α values.


b)


Measurement of


a


npn BJT in a particular circuit shows the base current to


be 14.46 μA, the emitter current to be 1.460 mA, an


d the base



emitter voltage


to be 0.7 V. For these conditions, calculate α, β, and


𝐼


𝑆


.


c)


Consider


a


npn transistor with


𝑉


𝐵𝐸


= 0.7 V at


𝐼


𝐶


= 1 mA. Find


𝑉


𝐵𝐸


at


𝐼


𝐶


=


0.1 mA


.



1
Expert's answer
2021-07-05T04:55:34-0400

The relationship between α and β is given by

"\alpha = \frac{\beta}{\beta +1}"

Part a

With β = 50

"\alpha = \frac{50}{50 +1} = 0.9804"

Part b

With β = 150

"\alpha = \frac{150}{150 +1} = 0.993"

α value ranges between 0.98 and0.99